the ball is tied on a string to whirl it having a radius of 50 cm at a velocity of 1.2 m per squared what will be the acceleration of the ball
1
Expert's answer
2021-10-28T08:52:06-0400

"a_n=v^2/r=1.2^2/0.5=2.88 \ (m/s^2)"